OpenStudy (anonymous):

A wooden pyramid, 6 inches tall, has a square base. A carpenter increases the dimensions of the wooden pyramid by a factor of 3 and makes a larger pyramid with the new dimensions. Describe in complete sentences the ratio of the volumes of the two pyramids

sam (.sam.):

\[V=\frac{1}{3}Bh\] ----------------------- Ratio \[V_1=3V_2\] \[\frac{1}{3}Bh_1=\cancel3(\frac{1}{\cancel3}Bh_2)\] \[\frac{1}{3}\cancel Bh_1=\cancel Bh_2\]

sam (.sam.):

\[\frac{1}{3}h_1=(6)\] \[h_1=18\]

OpenStudy (anonymous):

so what is the ratio of the two pyramids??

sam (.sam.):

Lets say B=5 \[V_1:V_2\] \[\frac{1}{3}(5)(6):\frac{1}{3}(5)(18)\] 10:30 1:3